# Persuasion in Networks: Can the Sender Do Better than Using Public Signals?

### v2

@@ 6 Conclusion
There has been a lot of discussion on exploiting network information for advertising and political campaigns. This paper studies whether and when using network signals outperforms public signals. In this paper, receivers have heterogeneous priors and pass on a signal to their neighbours if they are persuaded by the signal. This paper finds that risk netural and risk loving senders prefer public signals, if believers are more connected than sceptics. This holds even if the network exhibits homophily. The sender may prefer network signals if she is risk averse or scceptics are sufficiently more connected than believers.
The main intuition underlying results in this paper is that, the sender needs the signal to get viral in order to have an impact. However, it is difficult to control who can observe viral signals using seeding. The sender would like to send more information to sceptics than believers, but this is only possible when sceptics are more connected than believers.
In many situations, the campaigner has access to tools that can send public signals. The results in this paper imply that, in many cases, resources may be more efficiently spent on sending public signals, rather than on trying to understand network structures and picking the optimal seeds. In addition, persuasion on large networks is complicated, while public signals have been well studied. From these comparisons, we can use sender’s value from public signals to obtain what sender could get at most or at least using network signals. Even if the sender has no access to public signals, the results could be useful to judge whether a campaign using network signals will be profitable for the sender.
@@ A.1 Proof of main results
Using the lemmas 2, 3 and 4, I can now approximate probabilities that receivers observe an non-empty signal $s\neq\emptyset$. Let $C(.)$ and $\hat{C}(.)$ be functions over the set of seeds and networks defined in the following way
$C(z,g^{n})=\begin{cases}1&\text{ if }z\cup L_{1}\neq\emptyset\\ 0&\text{otherwise}\end{cases}\qquad\hat{C}(z,g^{n})=\begin{cases}1&\text{ if }z\cup\hat{L}_{1}\neq\emptyset\\ 0&\text{otherwise}\end{cases}$
@@
By lemmas 3 and 4, for any $\epsilon>0$, we can find $n$ large enough such that with arbitrarily high probability, the proportion of receivers with type $t$ and degree $d$ who observe $s$ conditional on seeds being on relevant giant component is within $[\zeta(t,d)+2\epsilon,\zeta(t,d)-2\epsilon]$ if $s$ persuades both types of receivers, and is within $[\hat{\zeta}(t,d)+2\epsilon,\hat{\zeta}(t,d)-2\epsilon]$ if $s$ persuades only type $h$ receivers. Result then follows from definition of functions $C(.)$ and $\hat{C}(.)$, and lemma 2. ∎
@@
Given such $1-\rho(\infty|l,1)$, if $\alpha=1$, the LHS of both equations must be larger than the RHS. The largest value of $\alpha$ is when $\alpha(1-\rho(\infty|l,1))=1$. WHen $\alpha$ takes this value, the LHS of the first equation is 1 and RHS is less than 1. Also, notice that the LHS of the first equation is linear in $\alpha$ while the RHS is convex. These facts imply that LHS of the first equation must always be larger than the RHS for all $\alpha\geq 1$. This concludes the proof. ∎
@@
Using this function, the next lemma formally presents (part of) the imitation argument illustrated in main text.
@@
Proof of Proposition 2 point 1:
For any $\beta^{\prime}>0$, we can find a $\beta>0$ such that $sup(v(x+8\beta)-v(x))<\beta^{\prime}$. For this $\beta$, we can find some $\overline{d}$ such that $\sum_{d>\overline{d}}(p_{d}^{h}+p_{d}^{l})<\beta$. Fix this $\overline{d}$. By Lemma 7, Lemma 8 and Remark 1, we can find some $n_{1}$ such that for all $n>n_{1}$, we have $\hat{V}(d,\Pi^{n})\leq V_{pub}^{*}+\beta^{\prime}$ for all $\Pi^{n}$ and the following conditions are satisfied with probabilty at least $(1-\beta)$:
@@
Since $\hat{V}(d,\Pi^{n}_{1})\leq V_{pub}^{*}+\beta^{\prime}$ and $\hat{V}^{l}(d_{min},\Pi_{1}^{n})\leq V_{pub}^{*}$ (by 2), the first point follows since $\beta$ and $\beta^{\prime}$ can be arbitrarily small. ∎
Proof of Proposition 2 point 2:
Let the sender use the following network signalling strategy for all $n$
It can be verified that for any $r>1$, expression A.8 is strictly larger than 0.5 if $\hat{\zeta}(d,l)=1$ (see e.g. Innocenti (2022) for a proof). We know that $\hat{\zeta}(d,l)$ goes to 1 as $d$ approaches infinity. Therefore, we can find some $\underline{d}$ large enouch such that for all larger $d$, their posterior convergse to some number larger than 0.5. Let the payoff function $v(.)$ be defined as follows, $v(x)=x$ for all $x\leq\underline{x}$ and is equal to $\underline{x}$ for all larger $x$, where $\underline{x}>\gamma_{h}$ and $\underline{x}<\gamma_{h}+\gamma_{l}\sum_{d\geq\underline{d}}p_{d}^{l}(1-\hat{\zeta}(d,l))$. For $n$ large enough, sender gets payoff $\underline{x}$ from the network strategy described above. However, with public signals, it’s impossible to always persuade some type $l$ receivers, so sender’s payoff must be less than $\underline{x}$. ∎
Proof of 2: Sender’s payoff from the strategy described in proof of Proposition 2 point 2 is at least $\frac{\overline{x}^{\beta}-1}{\beta}$, and from public signal is at most $\mu_{s}(\omega=0)(1-\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)})\frac{\gamma_{h}^{\beta}-1}{\beta}$. The former will be larger than latter for $\beta$ small enough. ∎
@@
Consider the following network signal strategy to be used for all $n$: $\pi(s|\omega=1)=1$, $\pi(s|\omega=0)=\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}$ and $\pi(\emptyset|\omega)=1-\pi(s|\omega)$. Seed is always chosen to be on $L_{1}$, the giant component. Type $h$ nodes with degree 00 always chooses action 1, and type $l$ receivers who observe signal $s$ chooses action 1. Sender’s payoff is at least
@@
A.2 Proof for Voting Game
@@
We know that the proportion of type $h$ receivers converge to $\gamma_{h}$, and $\sum_{d\geq d_{vote}}\frac{N_{l,d}}{n}$ converges in probability to $\sum_{d\geq d_{vote}}\gamma_{l}p_{d}^{l}$ (by remark 1). Therefore, for fixed $\beta$ satisfying conditions above and any $\overline{d}>d_{vote}$ (if $d_{vote}$ is infinity, then select $\overline{d}$ to be large enough such that $\gamma_{h}+\sum_{d>\overline{d}}\gamma_{l}p_{d}^{l}<\overline{x}$), we can find some $n_{1}$ such that for all $n>n_{1}$, with probability at least $1-\beta$, we have $g^{n}\in\mathbf{A}^{n}(\beta,\overline{d})$
@@
But this inequality is a contradiction with how $\beta$ is chosen, so $a^{*}(\emptyset,l,d;\Pi^{n})=0$ for all $d<min\{d_{vote},\overline{d}\}$. Therefore, an upper bound on sender’s payoff when signal $s$ does not persuade type $l$ receivers is $v(\frac{1}{n}(N_{h}+\sum_{d\geq min\{d_{vote},\overline{d}\}}N_{l,d}))$ by assuming that all type $h$ receivers take action 1, and all type $l$ receivers with degree higer than $min\{d_{vote},\overline{d}\}$ choose action 1. By remark 1, this converges in probability to $v(\gamma_{h}+\sum_{d\geq min\{d_{vote},\overline{d}\}}\gamma_{l}p_{d}^{l})=0$, since $\gamma_{h}+\sum_{d\geq min\{d_{vote},\overline{d}\}}\gamma_{l}p_{d}^{l}<\overline{x}$. Therefore, sender can only get 0 unless a signal that persuades both types is realised, which implies that $\limsup_{n}V^{*}_{net}(n)$ is weakly smaller than $\sum_{\omega}\mu_{s}(\omega)\sum_{s\in S_{good}}\pi(s|\omega)$ for $n$ large enough. Since $\sum_{s\in S_{good}}\pi(s|\omega=1)\leq 1$ and
$\sum_{s\in S_{good}}\pi(s|\omega=0)\leq\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}$, this upper bound is weakly smaller than $V_{pub}^{*}$ found above. This concludes proof for the necessary condition.
To show the sufficient condition, suppose that sender uses strategy $\Pi_{2}^{n}$ described as follows for all $n$,
@@
If the sufficient condition holds, we can find some $\overline{d}$ such that $\gamma_{h}+\sum_{d=d_{vote}+1}^{\overline{d}}\gamma_{l}p_{d}^{l}(1-\hat{\zeta}(l,d))>\overline{x}$. By the argument above, for $n\geq n_{2}$, the proportion of receivers choosing action 1 in any realization is at least $\frac{1}{n}(N_{h}+\sum_{d=d_{vote}+1}^{\overline{d}}(N_{l,d}-N_{l,d,\hat{L}_{1}}(g^{n})))$. By remark 1 and Lemma 4, it converges in probability to $\gamma_{h}+\sum_{d=d_{vote}+1}^{\overline{d}}\gamma_{l}p_{d}^{l}(1-\hat{\zeta}(l,d))>\overline{x}$, so sender’s payoff converges in probability to $v(1)$. This is strictly larger than $V_{pub}^{*}=\mu_{s}(\omega=1)v(1)+\mu_{s}(\omega=0)\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}v(1)$. ∎
A.3 Proofs for technical lemmas in section 3
@@
Conditional on signal $s$ and a realised network, the set of nodes being selected as seeds and thus receivers observing the signal is completely determined, and therefore does not depend on the state $\omega$. Therfore, we know that
$Pr(\text{observing }s|s\text{ is realised},\omega,d,t,g^{n})$
is independent of $\omega$. In addition, we know that conditional on $s$, the network formation process is independent of state, i.e. $Pr(\omega,g^{n},d,t|s\text{ is realised})=Pr(\omega|s\text{ is realised})Pr(g^{n},d,t|s\text{ is realised})$. This implies that $Pr(g^{n}|s\text{ is realised},\omega,d,t)$ is independent of $\omega$. Therefore, we have
$Pr(\text{observing }s|s\text{ is realised},\omega,d,t)=Pr(\text{observing }s|s\text{ is realised},d,t)$
@@
I’ll prove lemmas Lemma 3 and Lemma 2 together in the following proof.
@@
If it changes the component structure. It can only break some component into 2. If that component that’s broken is of size $m$, then it will change the number of nodes with the property described by at most $m+2$. If it’s of size more than $m$, it’s possible that it’s broken into two compoenents of size $m$. Therefore, it can change the number of nodes with the desired property by at most $22m+$ 2. ∎
@@
Conditional on having degree $d$ and type $t$ and $\lambda$ with $f_{t}(\lambda)>0$, the proportion of nodes on the $L_{1}$ converges in probability to $\rho(\infty|t,d)$, which is increasing in $d$.
@@
Conditional on having degree $d$ and type $t$, the proportion of nodes with a neighbour on the $\hat{L}_{1}$ converges in probability, and the limit is increasing in $d$
@@
Using the argument preceeding 1 again, the distribution over subnetwork among type $h$ nodes is the same as the distribution over networks after setting $\lambda_{i}=0$ for all type $l$ nodes. I use $\hat{\rho}(\infty)$ to denote the probability that the associated process never dies out, and similarly $\hat{\rho}(\infty|\cdot)$ to denote the conditional probabilities.
@@
The probability that a node of type $l$ and connectedness $\lambda$ is connected to a node of type $h$ and connectedness $\lambda^{\prime}$ is $\frac{\lambda\lambda^{\prime}q}{n}$. Given types and connectedness of a pair of nodes, whether they are connected are independent of the types of all other nodes, and formation of all other links. Therefore, conditional on the type $h$ subnetwork, a node of type $l$, connectedness $\lambda$ and $d_{l}$ type $l$ neighbours is connected to 00 nodes on $\hat{L}_{1}$ and $k$ type $h$ nodes not on $\hat{L}_{1}$ with probability
@@
$\frac{N_{l,d_{l},d_{h},\lambda}-N_{l,d_{l},d_{h},\lambda,\hat{L}_{1}}}{N_{l,d_{l},\lambda}}=PO(d_{h};\sum_{\lambda^{\prime}}\lambda\lambda^{\prime}q\gamma_{h}f_{h}(\lambda^{\prime}))Binom(d_{h};d_{h},\frac{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{h}(\lambda^{\prime})(1-\hat{\rho}(\infty|h,\lambda^{\prime}))}{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{h}(\lambda^{\prime})})$
@@
which converges to $PO(k;\sum_{\lambda^{\prime}}\lambda\lambda^{\prime}q\gamma_{h}f_{h}(\lambda^{\prime}))$. Therefore, $\frac{N_{l,d_{l},d_{h},\lambda}}{N_{l,d_{l},\lambda}}$ converges in probability to $PO(d_{h};\sum_{\lambda^{\prime}}\lambda\lambda^{\prime}q\gamma_{h}f_{h}(\lambda^{\prime}))$. Therefore,
$\frac{N_{l,d_{l},d_{h},\lambda,\hat{L}_{1}}}{N_{l,d_{l},d_{h},\lambda}}=1-(\frac{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{h}(\lambda^{\prime})(1-\hat{\rho}(\infty|h,\lambda^{\prime}))}{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{h}(\lambda^{\prime})})^{d_{h}}=\hat{\rho}(\infty|d_{h},h)$
where the last equality follows from 4 and constuction of the network formatio process associated with $\hat{\rho}(\cdot)$. This implies that $\frac{N_{l,d_{l},d_{h},\hat{L}_{1}}}{N_{l,d_{l},d_{h}}}=\hat{\rho}(\infty|d_{h},h)$, and

### v3

@@ 6 Conclusion
There has been a lot of discussion on exploiting social networks to spread information and persuade people in political and advertising campaigns. This paper studies whether and when network signals outperform public signals. This paper finds that a risk neutral or risk-loving sender prefers public signals if believers are more connected than sceptics. This holds even if the network exhibits homophily. The sender prefers network signals if sceptics are sufficiently more connected than believers. When believers are sufficiently connected, a very risk averse sender prefers network signals. If the sender’s payoff function is a threshold function, connectedness of believers is also important for the comparison.
The main intuition underlying results in this paper is that, the sender needs the signal to get viral in order to have an impact. However, it is difficult to control who can observe viral signals using seeding. The results specify how different networks spread viral signals, and whether the sender can exploit it to outperform public signals.
In many situations, the campaigner has access to tools that can send public signals. The results in this paper imply that, in many cases, resources may be more efficiently spent on sending public signals, rather than on trying to understand network structures and picking the optimal seeds. Conditions under which network signals perform better are also specified, and there is a stronger case to collect more information about the network under these conditions. In addition, persuasion on large networks is complicated, while public signals have been well studied. From these comparisons, we can use sender’s value from public signals to obtain what sender could get at most or at least using network signals. Even if the sender has no access to public signals, the results could be useful to judge whether a campaign using network signals will be profitable for the sender.
@@ A.1 Proof of main results
[…]
$C(z,g^{n})=\mathbf{1}(z\cup L_{1}(g^{n})\neq\emptyset)\qquad\hat{C}(z,g^{n})=\mathbf{1}(z\cup\hat{L}_{1}(g^{n})\neq\emptyset)$
@@
By lemmas 3 and 4, for any $\epsilon>0$, we can find $n$ large enough such that with arbitrarily high probability, the proportion of receivers with type $t$ and degree $d$ who observe $s$ conditional on seeds being on relevant giant component is within $\zeta(t,d)+2\epsilon,\zeta(t,d)-2\epsilon]$ if $s$ persuades both types of receivers, and is within $\hat{\zeta}(t,d)+2\epsilon,\hat{\zeta}(t,d)-2\epsilon]$ if $s$ persuades only type $h$ receivers. Result then follows from definition of functions $C(.)$ $\hat{C}(.), and$ lemma 2. ∎
@@
Given such $1-\rho(\infty|l,1)$, if $\alpha=1$, the LHS of both equations must be larger than the RHS. The largest value of $\alpha$ is when $\alpha(1-\rho(\infty|l,1))=1$. WHen $\alpha$ takes this value, the LHS of the first equation is 1 and RHS is less than 1. Also, notice that the LHS of the first equation is linear in $\alpha$ while the RHS is convex. These facts imply that LHS of the first equation must always be larger than the RHS for all $\alpha\geq 1$. This concludes the proof for Lemma 6. The analogous argument shows that $\alpha$ cannot be smaller than 1 when degree distribution and thus forward distributions are the same for both types. This shows Lemma 5. ∎
@@
Using this function, the next lemma formally presents (part of) the imitation argument.
@@
Proof of Proposition 2:
For any $\beta^{\prime}>0$, we can find a $\beta>0$ such that $sup(v(x+8\beta)-v(x))<\beta^{\prime}$. For this $\beta$, we can find some $\overline{d}$ such that $\sum_{d>\overline{d}}(p_{d}^{h}+p_{d}^{l})<\beta$. Fix this $\overline{d}$. By Lemma 7, Lemma 8 and 1, we can find some $n_{1}$ such that for all $n>n_{1}$, we have $\hat{V}(d,\Pi^{n})\leq V_{pub}^{*}+\beta^{\prime}$ for all $\Pi^{n}$ and the following conditions are satisfied with probabilty at least $(1-\beta)$:
@@
Since $\hat{V}(d,\Pi^{n}_{1})\leq V_{pub}^{*}+\beta^{\prime}$ and $\hat{V}^{l}(d_{min},\Pi_{1}^{n})\leq V_{pub}^{*}$ (by 3), the first point follows since $\beta$ and $\beta^{\prime}$ can be arbitrarily small. ∎
@@
Consider the following network signal strategy to be used for all $n$: $\pi(s|\omega=1)=1$, $\pi(s|\omega=0)=\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}$ and $\pi(\emptyset|\omega)=1-\pi(s|\omega)$. Seed is always chosen to be on $L_{1}$, the giant component. Type $h$ nodes with degree 0 always chooses action 1, and type $l$ receivers who observe signal $s$ chooses action 1. Sender’s payoff is at least
@@
Proof of Proposition 4:
Let the sender use the following network signal strategy for all $n$
It can be verified that for any $r>1$, expression A.8 is strictly larger than 0.5 if $\hat{\zeta}(d,l)=1$ (see e.g. Innocenti (2022) for a proof). We know that $\hat{\zeta}(d,l)$ goes to 1 as $d$ approaches infinity. Therefore, we can find some $\underline{d}$ large enouch such that for all larger $d$, their posterior convergse to some number larger than 0.5. Let the payoff function $v(.)$ be defined as follows, $v(x)=x$ for all $x\leq\underline{x}$ and is equal to $\underline{x}$ for all larger $x$, where $\underline{x}>\gamma_{h}$ and $\underline{x}<\gamma_{h}+\gamma_{l}\sum_{d\geq\underline{d}}p_{d}^{l}(1-\hat{\zeta}(d,l))$. For $n$ large enough, sender gets payoff $\underline{x}$ from the network strategy described above. With public signals, it’s impossible to always persuade some type $l$ receivers, so sender’s payoff must be less than $\underline{x}$.
If the sender has the constant relative risk aversion utility described in proposition, the sender’s payoff from the strategy described above [^16] Note that the sender gets slightly less payoff from $s^{\prime}$ compared to $\emptyset$. Nonetheless, when $b$ gets large, the payoff function will be arbitrarily close to flat beyond $\overline{x}$ defined in the last paragraph. Therefore, we can find some function that arbitrarily close to the constant relative risk aversion payoff function, where the sender gets weakly lower payoff from $\emptyset$ than $s^{\prime}$. is at least $\frac{\overline{x}^{b}-1}{b}$, and from public signal is at most
$\mu_{s}(\omega=0)(1-\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)})\frac{\gamma_{h}^{b}-1}{b}$. The former will be larger for $b$ small enough. ∎
Proof for Voting Game:
@@
We know that the proportion of type $h$ receivers converge to $\gamma_{h}$, and $\sum_{d\geq d_{vote}}\frac{N_{l,d}}{n}$ converges in probability to $\sum_{d\geq d_{vote}}\gamma_{l}p_{d}^{l}$ (by 1). Therefore, for fixed $\beta$ satisfying conditions above and any $\overline{d}>d_{vote}$ (if $d_{vote}$ is infinity, then select $\overline{d}$ to be large enough such that $\gamma_{h}+\sum_{d>\overline{d}}\gamma_{l}p_{d}^{l}<\overline{x}$), we can find some $n_{1}$ such that for all $n>n_{1}$, with probability at least $1-\beta$, we have $g^{n}\in\mathbf{A}^{n}(\beta,\overline{d})$
@@
But this inequality is a contradiction with how $\beta$ is chosen, so $a^{*}(\emptyset,l,d;\Pi^{n})=0$ for all $d<min\{d_{vote},\overline{d}\}$. Therefore, an upper bound on sender’s payoff when signal $s$ does not persuade type $l$ receivers is $v(\frac{1}{n}(N_{h}+\sum_{d\geq min\{d_{vote},\overline{d}\}}N_{l,d}))$ by assuming that all type $h$ receivers take action 1, and all type $l$ receivers with degree higer than $min\{d_{vote},\overline{d}\}$ choose action 1. By 1, this converges in probability to $v(\gamma_{h}+\sum_{d\geq min\{d_{vote},\overline{d}\}}\gamma_{l}p_{d}^{l})=0$, since $\gamma_{h}+\sum_{d\geq min\{d_{vote},\overline{d}\}}\gamma_{l}p_{d}^{l}<\overline{x}$. Therefore, sender can only get 0 unless a signal that persuades both types is realised, which implies that $\limsup_{n}V^{*}_{net}(n)$ is weakly smaller than $\sum_{\omega}\mu_{s}(\omega)\sum_{s\in S_{good}}\pi(s|\omega)$ for $n$ large enough. Since $\sum_{s\in S_{good}}\pi(s|\omega=1)\leq 1$ and $\sum_{s\in S_{good}}\pi(s|\omega=0)\leq\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}$, this upper bound is weakly smaller than $V_{pub}^{*}$ found above. This concludes proof for the necessary condition.
To show the sufficient condition, suppose that sender uses strategy $\Pi_{2}^{n}$ below for all $n$,
@@
If the sufficient condition holds, we can find some $\overline{d}$ such that $\gamma_{h}+\sum_{d=d_{vote}+1}^{\overline{d}}\gamma_{l}p_{d}^{l}(1-\hat{\zeta}(l,d))>\overline{x}$. By the argument above, for $n\geq n_{2}$, the proportion of receivers choosing action 1 in any realization is at least $\frac{1}{n}(N_{h}+\sum_{d=d_{vote}+1}^{\overline{d}}(N_{l,d}-N_{l,d,\hat{L}_{1}}(g^{n})))$. By 1 and Lemma 4, it converges in probability to $\gamma_{h}+\sum_{d=d_{vote}+1}^{\overline{d}}\gamma_{l}p_{d}^{l}(1-\hat{\zeta}(l,d))>\overline{x}$, so sender’s payoff converges in probability to $v(1)$. This is strictly larger than $V_{pub}^{*}=\mu_{s}(\omega=1)v(1)+\mu_{s}(\omega=0)\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}v(1)$. ∎
A.2 Proofs for technical lemmas in section 3
@@
[…]
@@
Proof of 2: It directly follows from Lemma 1 and the discussion after it. Lemma 1 pins down people’s behaviour for each $s\in S$, and thus the group who observe $s$ given network and seeds. These in turn pin down the probabilities that receivers observe each $s$, and thus also $\emptyset$ which is observed with complementary probability. From this, we can get receivers’ behaviour upon observing $\emptyset$. Given a signal and network, all these together determine each receiver’s observation and their behaviour, and thus the sender’s payoff. ∎
@@
I’$l$ prove lemmas Lemma 3 and Lemma 2 together in the following proof.
@@
If it changes the component structure. It can only break some component into 2. If that component that’s broken is of size $m$, then it will change the number of nodes with the property described by at most $m+2$. If it’s of size more than $m$, it’s possible that it’s broken into two compoenents of size $m$. Therefore, it can change the number of nodes with the desired property by at most $2m+2$. ∎
@@
Among receivers having degree $d$, type $t$ and connectedness $\lambda$ with $f_{t}(\lambda)>0$, the proportion of those on the $L_{1}$ converges in probability to $\rho(\infty|t,d)$, which is increasing in $d$.
@@
Among receivers having degree $d$ and type $t$, the proportion of those with a neighbour on the $\hat{L}_{1}$ converges in probability, and the limit is increasing in $d$
@@
[…]
@@
The probability that a node of type $l$ and connectedness $\lambda$ is connected to a node of type $h$ and connectedness $\lambda^{\prime}$ is $\frac{\lambda\lambda^{\prime}q}{n}$. Given types and connectedness of a pair of nodes, whether they are connected are independent of the types of all other nodes, and formation of all other links. Therefore, conditional on the type $h$ subnetwork, a node of type $l$, connectedness $\lambda$ and $d_{l}$ type $l$ neighbours is connected to 0 nodes on $\hat{L}_{1}$ and $k$ type $h$ nodes not on $\hat{L}_{1}$ with probability
@@
$\frac{N_{l,d_{l},d_{h},\lambda}-N_{l,d_{l},d_{h},\lambda,\hat{L}_{1}}}{N_{l,d_{l},\lambda}}\rightarrow_{p}PO(d_{h};\sum_{\lambda^{\prime}}\lambda\lambda^{\prime}q\gamma_{h}f_{h}(\lambda^{\prime}))Binom(d_{h};d_{h},\frac{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{h}(\lambda^{\prime})(1-\hat{\rho}(\infty|h,\lambda^{\prime}))}{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{h}(\lambda^{\prime})})$
@@
which converges to $PO(k;\sum_{\lambda^{\prime}}\lambda\lambda^{\prime}q\gamma_{h}f_{h}(\lambda^{\prime}))$.
Therefore, $\frac{N_{l,d_{l},d_{h},\lambda}}{N_{l,d_{l},\lambda}}$ converges in probability to $PO(d_{h};\sum_{\lambda^{\prime}}\lambda\lambda^{\prime}q\gamma_{h}f_{h}(\lambda^{\prime}))$, and
$\frac{N_{l,d_{l},d_{h},\lambda,\hat{L}_{1}}}{N_{l,d_{l},d_{h},\lambda}}\rightarrow_{p}1-(\frac{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{h}(\lambda^{\prime})(1-\hat{\rho}(\infty|h,\lambda^{\prime}))}{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{h}(\lambda^{\prime})})^{d_{h}}=\hat{\rho}(\infty|d_{h},h)$
[…]